OpenStudy (anonymous):

sketch the domain of f f(x,y)= arcsin(2x-y)/(x+y)

OpenStudy (anonymous):

let me rewrite the equation \[f(x,y)=\arcsin \frac{ 2x-y }{ x+y }\]

OpenStudy (anonymous):

@SolomonZelman help please

OpenStudy (michele_laino):

I think that the domain of the function above, is given by the solution of these inequalities: \[ - 1 \leqslant \frac{{2x - y}}{{x + y}} \leqslant 1,\quad x + y \ne 0\] wherein the first inequality, comes from the natural limits of \(\sin \) function whereas the second inequality, comes from the impossibility to divide by zero

OpenStudy (michele_laino):

so, first, we have to subtract the line \(y=-x\) from the \((x,y)-\)plane: 

OpenStudy (anonymous):

@Michele_Laino does the line of 2x-y need to draw too?

OpenStudy (michele_laino):

no, I don't think. Here we have to solve these two inequalities: \[\frac{{2x - y}}{{x + y}} \leqslant 1\] and: \[\frac{{2x - y}}{{x + y}} \geqslant - 1\]

OpenStudy (michele_laino):

the first inequality, is equivalent to these two systems of inequalities: \[\left\{ \begin{gathered} 2x - y \leqslant x + y \hfill \\ x + y > 0 \hfill \\ \end{gathered} \right. \cup \left\{ {\begin{array}{*{20}{c}} {2x - y \geqslant x + y} \\ {x + y < 0} \end{array}} \right.\]

OpenStudy (anonymous):

is this equation valid? 2x-y=0

OpenStudy (michele_laino):

the solutions of those systems are: 

OpenStudy (anonymous):

how can i know which part to shade??

OpenStudy (michele_laino):

since you have to solve the two systems above

OpenStudy (michele_laino):

for example, I solve the first system: \[\left\{ \begin{gathered} 2x - y \leqslant x + y \hfill \\ x + y > 0 \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

here is the next step: \[\left\{ \begin{gathered} 2x - y \leqslant x + y \hfill \\ x + y > 0 \hfill \\ \end{gathered} \right. \Rightarrow \left\{ \begin{gathered} x \leqslant 2y \hfill \\ y > - x \hfill \\ \end{gathered} \right. \Rightarrow \left\{ \begin{gathered} y \geqslant x/2 \hfill \\ y > - x \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

the first inequality is represented by all points \((x,y)\) such that \(y \geqslant x/2\), whereas the second inequality is represented by all points \((x,y)\) such that \(y>-x\). The solution is given by the intersection between both regions

OpenStudy (anonymous):

ohh.. i get it, thank you so much

OpenStudy (michele_laino):

ok! Now we have to solve the second system: \[\left\{ {\begin{array}{*{20}{c}} {2x - y \geqslant x + y} \\ {x + y < 0} \end{array}} \right.\]

OpenStudy (michele_laino):

I got this: \[\left\{ {\begin{array}{*{20}{c}} {2x - y \geqslant x + y} \\ {x + y < 0} \end{array}} \right. \Rightarrow \left\{ \begin{gathered} x \geqslant 2y \hfill \\ y < - x \hfill \\ \end{gathered} \right. \Rightarrow \left\{ \begin{gathered} y \leqslant x/2 \hfill \\ y < - x \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

he first inequality selects all points (x,y) such that \(y \leqslant x/2\), whereas the second inequality is represented by all points \((x,y)\) such that \(y<−x\). The solution is given by the intersection between both regions the solution of the first inequlity, is given by the union of the two solution of those two system. Such solution is represented in my drawing above

OpenStudy (michele_laino):

oops...the* first...

OpenStudy (michele_laino):

another typo: inequality*

OpenStudy (michele_laino):

now, in order to get the requested domain, you have to add, using the same procedure above, the solution of this second inequality: \[\frac{{2x - y}}{{x + y}} \geqslant - 1\]

OpenStudy (anonymous):

i dont get this part, why \[2x-y \] become\[y \le x/2\] shouldnt it be \[x \le y/2\]

OpenStudy (michele_laino):

if we start from this inequality: \[2x - y \leqslant x + y\] then I add \(y\) to both sides, so I can write: \[\begin{gathered} 2x - y + y \leqslant x + y + y \hfill \\ \hfill \\ 2x \leqslant x + 2y \hfill \\ \end{gathered} \] finally, I subtract \(x\) from both sides: :\[\begin{gathered} 2x - x \leqslant x + 2y - x \hfill \\ \hfill \\ x \leqslant 2y \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

with the same procedure, if I start with this inequality: \[{2x - y \geqslant x + y}\] I get: \[{x \geqslant 2 y}\]

OpenStudy (anonymous):

ohhh.. i see my mistake now

OpenStudy (michele_laino):

next step: the solution of the subsequent inequality: \[\frac{{2x - y}}{{x + y}} \geqslant - 1\] is given by the union of the solutions of these two systems of inequalities: \[\left\{ \begin{gathered} 2x - y \geqslant - x - y \hfill \\ x + y > 0 \hfill \\ \end{gathered} \right. \cup \left\{ {\begin{array}{*{20}{c}} {2x - y \leqslant - x - y} \\ {x + y < 0} \end{array}} \right.\]

OpenStudy (michele_laino):

please try to solve such systems of inequalities

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

we need to do both in order to obtain the graph?

OpenStudy (michele_laino):

yes! That's right!

OpenStudy (anonymous):

ok, thanks for the help :)